INFRARED SPECTRA OF X^{−}^{.}H_{2}O (X=F,Cl,Br) IN THE 600-800 cm^{−}^{1} REGION: UNDERSTANDING THE THREE-DIMENSIONAL CONFINEMENT OF THE SHARED PROTON

\underline{JOSEPH R. ROSCIOLI}, ERIC G. DIKEN, MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, PO Box 208107, New Haven, CT 06520.

Recent advances in nonlinear mixing technology have allowed us to obtain infrared spectra of the X^{−}^{.}H_{2}O (X=F,Cl,Br) complexes in the 600-1800 cm^{−}^{1} (16.7-5.6 μm) region. While transitions in this regime are typically associated with the three-dimensional confinement of the shared proton, unexpected spectral activity is observed. The out-of-plane vibrational overtone of the Cl^{−}^{.}H_{2}O complex exhibits anomalous intensity, which is explained by a curvilinear motion of the proton. The F^{−}^{.}H_{2}O spectrum shows evidence for Fermi interactions between several energy levels. Analysis of the spectra in the context of the potential surface probed by the shared proton is discussed.